Consider the UNILUS economy in the long-run described by the following equations:

Y = C + I + G, Y = 5,000, G = 1,000, T = 1,000, C = 250 + 0.75(Y − T), I = 1,000 − 50r. C=100+b(Y-50-0.25Y);


1
Expert's answer
2021-09-14T07:32:40-0400

"Y = C +I + G"


"C = 250 + 0.75(5000-1000)"


"C = 3000"


"5000 = 3000+ (1000-50r) + 1000"


"5000 = 4000-50r"


"5000-4000 = 50r"


"1000 = 50r"

"r = \frac{1000}{50} = 20"


"I = 1000-50r"

"= 1000-50(20) = 0"


"C = 100+b(Y0.5 -0.25Y)"

"3000 = 100+ b((0.5 \u00d7 5000) - (0.25\u00d75000))""3000 = 100 + b(2500-1250)"


"3000= 100 + b(1250)"


"3000-100 = 1250b"


"2900= 1250b"


"b= \frac{2900}{1250}"


"b= 2.32"